# ViCor: Bridging Visual Understanding and Commonsense Reasoning with Large Language Models

### v1

@@ 6.1 Ablation Study
On BLIP2-Pretrain, using clues for image-text alignment is not as good as using the transferred declarative sentences. This validates that BLIP2 can already understand high-level concepts well. However, introducing visual factors and observations as extra context improves performance on LLM reasoning, especially when the LLM is not confident about its initial judgment, i.e., initial provided visual information (caption) is insufficient. In this case, the performance of LLM reasoning (‘Cap + Clue’ in Table 1) is comparable with pre-trained BLIP2. This is because compared with BLIP2, LLM also takes caption as input and can implicitly combine the information from both caption and the visual clues.
For VCI problems, visual factors and visual clue generations help both reasoning paradigms. First, the significant improvement in the BLIP2-Pretrain paradigm validates that (1) pre-trained BLIP2 cannot well-align statements that goes beyond literal visual content, requiring commonsense inference; (2) LLM can reason about the visual factors that may contribute to supporting candidate commonsense inferences, and guide the VLM to focus on relevant factors accordingly.
Second, on LLM reasoning paradigm, the improvement of visual clue information is more significant than on the VCU problems with 9.5% vs. 7.4%. This shows that VCI requires the perception of subtle details of the scene compared to VCU problems to make sense of the current situation, which is harder to capture in a general caption. Third, visual clues reasoned by LLM are better than VQA as the visual information provider. There are mainly two reasons. First, the pre-trained VLM sometimes could not understand or correctly answer the question due to the lack of language alignment. Second, the VQA model lacks the main question as the context and may not get the intention of the visual factor. Therefore, it may produce irrelevant answers. We provide the examples to further illustrate these in Section 6.3.
How does confidence affect the reasoning process?
From Table. 1 we can observe that when the LLM is confident about its initial reasoning, the performance is better than BLIP2-Pretrain on both VCU and VCI problems. Second, combining more visual observations from VQA clue or LLM clue in these examples hurts the performance most of the time. Therefore, including more information is unnecessary in this case.
@@ 6.2 Main Results
The results on VCR dataset are in Table 3. Our method achieves the best result compared with other methods without supervised training. Specifically, our method outperforms the IdealGPT (You et al., 2023) since it is able to leverage the visual recognition and understanding abilities of VLMs more effectively by considering the types and definitions of problems. However, we notice that there is still a significant gap between ICL methods and methods with supervised training. This could be due to the loss of information in approximating the naming and labeling of the persons mentioned in Section 5.1.
@@ 6.3 Qualitative Examples
7 Conclusion and Limitations
In this work, we study the problem of visual commonsense reasoning (VCR) based on the capabilities of pre-trained vision-language models and large-language models and define two sub-problems – visual commonsense understanding (VCU) and visual commonsense inference (VCI). Based on this, we propose the ViCor framework that efficiently uses visual recognition and understanding capabilities of VLMs and commonsense reasoning capabilities of LLMs to overcome the challenges in VCR. The experiment results validate our analysis of VCR problems and the effectiveness of our framework. Currently, our method lags behind best performing methods in the field which are based supervised fine-tuning. Also, text is the only communication medium between LLMs and VLMs. The loss of visual details caused by captions may be hindering on certain scenarios. Future work could explore fine-tuning approach with alternative mediums such as visual embeddings.

### v2

@@ 6.1 Ablation Study
On BLIP2-Pretrain, using clues for image-text alignment is not better than using directly transferred declarative sentences. This validates that BLIP2 can already align visual features with different concepts well. However, introducing visual factors and observations as extra context improves performance on LLM reasoning, especially when the LLM is not confident about its initial judgment with only caption as context. In this case, the performance of LLM reasoning (‘Cap + Clue’ in Table 1) is comparable with pre-trained BLIP2.
For VCI problems, visual factors and visual clue generations help both reasoning paradigms. First, the improvement in the BLIP2-Pretrain paradigm validates that (1) pre-trained BLIP2 cannot well-align statements that go beyond literal visual content, requiring commonsense inference; (2) LLM can reason about the visual factors that may contribute to supporting candidate commonsense inferences, and guide the VLM to focus on relevant factors accordingly. Second, the improvement in the LLM reasoning paradigm shows that LLM clues successfully provide subtle details of the scene that are crucial for solving the problem. Third, visual clues reasoned by LLM are better than VQA as the visual information provider. There are mainly two reasons. First, the pre-trained VLM sometimes could not understand or correctly answer the question due to the lack of language alignment. Second, the VQA model lacks the main question as the context and may not get the intention of the visual factor. Therefore, it may produce irrelevant answers. We provide examples to further illustrate these in Section 6.3.
How to determine the reasoning process based on confidence and problem category for better LLM and VLM collaboration? When deciding the reasoning process, we need to consider both the performance and efficiency, evaluating by the number of LLM calls. From Table. 1, we can observe that when the LLM is confident about its initial reasoning, the performance is the best or almost the best on both VCU and VCI problems. Therefore, using LLM+caption is the best choice. When the LLM is not confident about its initial reasoning on VCI problems, LLM+Caption+LLM clue significantly outperforms other decision paradigms. On VCU problems, we can observe that the performance of BLIP2 is similar to LLM+Caption+LLM clue. However, the LLM+Caption+LLM clue requires five LLM calls, which is three times more than using BLIP2. Therefore, using BLIP2-ITA is the best choice in this case.
@@ 6.2 Main Results
The results on VCR dataset are in Table 3. Our method achieves the best result compared with other training-free methods. Specifically, our method outperforms IdealGPT (You et al., 2023) since it is able to leverage the visual understanding abilities of VLMs more effectively by considering the types and definitions of problems. However, we notice that there is still a significant gap between ICL methods and methods with supervised training. This could be due to the loss of information in approximating the naming and labeling of the persons mentioned in Section 5.1.
@@ 6.3 Qualitative Examples
7 Conclusion
In this work, we study the collaboration of pre-trained vision-language models and large-language models on a complex problem – visual commonsense reasoning (VCR). We analyze and validate the distinct advantages of LLMs and VLMs by testing them on two different types of VCR problems. Based on this, we propose the ViCor framework that efficiently uses the visual understanding capabilities of VLMs and commonsense reasoning capabilities of LLMs to overcome the challenges in VCR. The experiment results validate the effectiveness of our framework. We believe our study can provide insights into the roles and the collaboration of LLMs and VLMs in vision and language problems.
8 Limitation and Potential Risk
In our framework, we use text as the communication medium between LLMs and VLMs. The loss of visual details caused by captions may be hindering certain scenarios, and thus, our method lags behind supervised best-performing methods. Future work could explore incorporating our designs into an end-to-end fine-tuning approach. Large language models (LLMs) are a core component of our framework. Therefore, our method may inherit the potential risks from LLMs, such as hallucination and potentially offensive language.
Appendix A Additional Results
A.1 Results on More LLM Decoding Configurations
| Decision Model | Decoding Config | VCU | VCI | ||
| Conf | !Conf | Conf | !Conf | ||
| LLM + Caption + LLM Clue | Orig | 72.9 | 58.1 | 57.1 | 52.9 |
| Temp 0.1 | 72.4 | 58.8 | 57.1 | 53.9 | |
| Temp 0.2 | 72.4 | 58.5 | 61.2 | 52.2 | |
| ICL examples | 74.7 | 56.7 | 59.2 | 54.2 | |
| Num. of Examples | 170 | 1779 | 49 | 1002 | |
To validate the robustness of our method, we ran the experiments on the VCR dataset with more decoding configurations using LLM + Caption + LLM Clue decision branch. Specifically, we ran on two more LLM decoding temperatures 0.1 and 0.2, and used different in-context examples for the prompt in Fig.3 (right) to guide the LLM to think about observations for visual factors based on candidate choices. From the results in Table. 4, we can observe that different decoding configurations influence the results by a small margin and do not affect the main conclusions.
A.2 Results on OKVQA dataset
We adapt our method and baselines to OKVQA Marino et al. (2019) dataset. We use the full validation set, which contains 5046 examples. The results are in Table. 5. We use GPT-3.5-Turbo for LLM modules. Since OKVQA is an open-ended dataset, we use the Caption+VQA clue version of our method in Figure 1 to tackle unconfident VCI problems. As shown above, our framework can still leverage the advantage of both VLMs and LLMs to achieve better results thanks to the better collaboration between them, e.g., VLMs utilization based on problem classification and active visual information acquisition.
| Method | Accuracy |
| LLM+Caption | 34.6 |
| BLIP2-T5XL | 36.2 |
| ViCor (ours) | 38.7 |